40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Dec 1 02:39:00 2010, Suresh, Update, SUS, Installation of ETMU07 as ETMX 
    Reply  Thu Dec 2 02:50:12 2010, kiwamu, Update, SUS, ETMX installed 
Message ID: 4002     Entry time: Wed Dec 1 02:39:00 2010     Reply to this: 4006
Author: Suresh 
Type: Update 
Category: SUS 
Subject: Installation of ETMU07 as ETMX 

[Kiwamu, Jenne, Koji, Suresh]

The following steps in this process were completed.

1)  Secured the current ETMX (Old ETMY) with the earth quake stops.

2) Removed the OSEMs and noted the Sl no. of each and its position

3) Placed four clamps to mark the location of the current ETMX tower (Old ETMY's position on the table)

4) Moved the ETMX (Old ETMY) tower to the clean table flow bench.  In the process the tower had to be tilted during removal because it was too tall to pass upright through the vacuum chamber port.  It was scary but nothing went wrong.

5) Koji calculated the location of the new ETMX and told us that it should be placed on the north end of the table.

6) Moved the OSEM cables, counter balancing weights and  the 'chopper' out of the way.  Had to move some of the clamps securing the cables.

7) Moved the ETMU07 tower from the clean room to the ETMX table

8) Positioned the OSEMs as they were placed in the earlier tower and adjusted their position to the middle of the range of their shadow sensors.  The four OSEMs on the face did not give us any trouble and were positioned as required.  But the side OSEM could not be put in place.  The magnet on the left side, which we are constrained to use since the tower is not designed to hold an OSEM on the right side, seems a little too low (by about a mm) and does not interrupt the light beam in the shadow sensor.  The possible causes are

   a) the optic is rotated.  To check this we need to take the tower back to the clean room and check the location of the optic with the traveling microscope.  If indeed it is rotated, this is easy to correct.

   b) the magnet is not located at the correct place on the optic.  This can also be checked on the clean room optical bench but the solution available immediately is to hold the OSEM askew to accommodate the magnet location.  If time permits the magnet position can be corrected.

We have postponed the testing of the ETMU07 tower to 1st of Nov Dec.

 

ELOG V3.1.3-